# Low-Rank Bandits via Tight Two-to-Infinity Singular Subspace Recovery

### v1

@@ D.3 Additional Lemmas
The control of the lower bound identified above is tight when the context distribution and the behavior policy are homegeneous and $\kappa=\Theta\left(1\right)$, since in this case, both sides of the inequality have the same scaling. In particular, the relaxed lower bound scales as the lower bound of Theorem 6.3.
@@
We are now ready to derive the worst-case lower bound of Proposition 6.4. To that end, we will show that $K_{M,\pi}:=\frac{1}{\sigma_{1}\omega_{\max}}\max\left(\sum_{i}\|Q^{i}_{\pi}\|_{2}^{2},\sum_{j}\|P^{j}_{\pi}\|_{2}^{2}\right)$ is of order $\frac{m+n}{\omega_{\max}mn}$ in the worst case (i.e., for a choice of homogeneous reward matrix and target policy that maximizes the lower bound). We use the notation $\pi_{j}:=\sum_{i=1}^{m}\rho_{i}\pi\left(j|i\right)$. We note that the third bullet point of the following proposition corresponds to the statement of Proposition 6.4.
@@
    $T\gtrsim\frac{\left(m+n\right)}{\omega_{\max}mn\varepsilon^{2}}\operatorname{kl}\left(\delta,1-\delta\right).$
@@
Since $P=U\sqrt{\Sigma}$ and $Q=V\sqrt{\Sigma}$, we have $\max_{i}\|P_{i}\|_{2}^{2}=\max_{i}\sum_{j=1}^{r}\sigma_{j}U_{x,j}^{2}\leq\sigma_{1}\|U\|_{2\to\infty}^{2}$ and similarly, $\max_{j}\|Q_{j}\|_{2}^{2}\leq\sigma_{1}\|V\|_{2\to\infty}^{2}.$ This finally ensures
@@
$T\geq\frac{\operatorname{kl}\left(\delta,1-\delta\right)}{2\left(m\vee n\right)\omega_{\max}\varepsilon^{2}}.$
@@ Appendix F Proofs of results from Section 7: Best Policy Identification
In this section, we denote by $\Pi$ the set of deterministic policies (in other words, the set of functions from $[m]$ to $[n]$). To ease the notation, we also denote by $\hat{v}^{\pi}_{\text{\tt RS}}$ the RS-PE estimator of $v^{\pi}$ for a given policy $\pi$, and by $\hat{\pi}_{\text{\tt RS}}$ the policy learned by RS-BPI.
@@
$T\gtrsim\frac{\mu^{2}r\left(m+n\right)}{\omega_{\min}mn\,\varepsilon^{2}}\log\left(\frac{m+n}{\delta}\right)+\frac{K_{2}}{\varepsilon}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$
where $K_{2}$ has been defined in (37).
@@
By replacing $\phi_{\pi}$ with $\phi_{i,j}$ in the error bound analysis performed in Appendix D, we directly obtain that for every (context,arm) pair $(i,j)$, with probability larger than $1-\frac{\delta}{mn}$,
@@
$T\gtrsim\frac{\mu^{2}r\left(m+n\right)}{\omega_{\min}mn\,\varepsilon^{2}}\log\left(\frac{emn}{\delta}\right)+\frac{K_{2}}{\varepsilon}\log^{3}\left(\frac{mn(m+n)T}{\delta}\right).$
@@
with probability larger than $1-\delta.$
@@
$v^{*}-v^{\hat{\pi}_{\text{\tt RS}}}\leq\varepsilon\text{ for }T\gtrsim\frac{\mu^{2}r\left(m+n\right)}{\varepsilon^{2}}\log\left(\frac{m+n}{\delta}\right)+\frac{K_{2}}{\varepsilon}\log^{3}\left(\frac{(m+n)T}{\delta}\right).$
We finally note that from (37), $K_{2}=\Theta\left(mn\right)$ when $M$ is also homogeneous.
We can of course derive the corresponding result for SBPI in the same way. More specifically, we obtain that SBPI is $\left(\varepsilon,\delta\right)$-PAC as soon as
$T\gtrsim\frac{\mu^{11}\kappa^{4}r^{3}\|M\|_{\max}^{2}(m+n)^{3}}{(m\wedge n)^{2}\varepsilon^{2}}\log^{3}\!\left(\frac{(m+n)T}{\delta}\right)\!$
by combining the max-norm guarantee on $\hat{M}$ of Proposition D.1 with the previous arguments. ∎
@@ F.2 General context distribution case
We now explain how to drop the homogeneous context assumption in Theorem 7.1 while retaining the optimal scaling up to logarithmic factors with the technique highlighted in (Lee et al., 2023). The idea is to split the set of contexts into subsets on which the context distribution is homogeneous and to apply RS-BPI or SBPI on each subset. Let $L=\left\lceil\log_{2}\left(m/\varepsilon\right)\right\rceil$ and define $\mathcal{C}_{l}=\left\{i\in[m],2^{-l-1}<\rho_{i}\leq 2^{-l}\right\}$ for $10\leq l\leq L-$ 1, and $\mathcal{C}_{L}=\left\{i\in[m],\rho_{i}\leq 2^{-L}\right\}$. On each $\mathcal{C}_{l}$ for $10\leq l\leq L-$ 1, the distribution defined by the normalized context weights is homogeneous. On $\mathcal{C}_{L}$, the context weights are small enough to not meaningfully contribute to the BPI error. We summarize below the algorithm structure for RS-BPI, and we recall that $\alpha$ is the proportion of samples used in the first phase of the algorithm. Of course, a SBPI version of this algorithm can be constructed in the same way.
@@
  Input: Budget of rounds $T$, data splitting parameter $\alpha$, context distribution $\rho$, regularization parameter $\tau$, confidence level $\delta$
@@
  Output: $\hat{\pi}_{\text{\tt RS}}$ such that $\hat{\pi}_{\text{\tt RS}}=\hat{\pi}^{l}_{\text{\tt RS}}$ on $\mathcal{C}_{l}$ for $10\leq l\leq L-$ 1 and arbitrary on $\mathcal{C}_{L}$.
@@
Let $\varepsilon,\delta\in(0,1)$, and define $\rho\left(\mathcal{C}_{l}\right):=\sum_{i\in\mathcal{C}_{l}}\rho_{i}$. With the choices $T_{1}=\lfloor\alpha T\rfloor$ and $\tau\leq r(m\wedge n)^{-1}(\rho_{\min}/\rho_{\max})\log(16d/\delta)$, G-RS-BPI is $(\varepsilon^{\prime},\delta)$-PAC as soon as
$T\gtrsim\frac{\mu^{2}r\left(m+n\right)}{\varepsilon^{2}}\log\left(\frac{L\left(m+n\right)}{\delta}\right)+\frac{K_{2}}{\varepsilon}\log^{3}\left(\frac{L(m+n)T}{\delta}\right)\!,$
where $K_{2}$ has been defined in (37) and $\varepsilon^{\prime}=\varepsilon\left(L+2\|M\|_{\max}\right)$.
@@
Let $10\leq l\leq L-$ 1 and $\delta_{l}=\delta/L$. From Lemma 9 in (Lee et al., 2023), we know that with high probability, at least
$\frac{T\rho\left(\mathcal{C}_{l}\right)}{2}\gtrsim\frac{\mu^{2}r\left(m+n\right)}{2\varepsilon_{l}^{2}}\log\left(\frac{m+n}{\delta_{l}}\right)+\frac{K_{2}}{\varepsilon_{l}}\log^{3}\left(\frac{(m+n)T}{\delta_{l}}\right)$
out of the $T$ samples correspond to a context in $\mathcal{C}_{l}$ for $\varepsilon_{l}:=\varepsilon/\rho\left(\mathcal{C}_{l}\right).$ Applying Algorithm 1 on $\mathcal{C}_{l}$ yields a policy $\hat{\pi}^{l}_{\text{\tt RS}}$ such that
$\sum_{i\in\mathcal{C}_{l}}\frac{\rho_{i}}{\rho\left(\mathcal{C}_{l}\right)}\left(M_{i,\hat{\pi}^{l}_{\text{\tt RS}}(i)}-M_{i,\pi^{*}(i)}\right)\leq\frac{\varepsilon}{\rho\left(\mathcal{C}_{l}\right)}$
@@
$\sum_{i=1}^{m}\rho_{i}\left(M_{i,\hat{\pi}_{\text{\tt RS}}(i)}-M_{i,\pi^{*}(i)}\right)$ $=\sum_{l=0}^{L-1}\sum_{i\in\mathcal{C}_{l}}\rho_{i}\left(M_{i,\hat{\pi}^{l}_{\text{\tt RS}}(i)}-M_{i,\pi^{*}(i)}\right)+\sum_{i\in\mathcal{C}_{L}}\rho_{i}\left(M_{i,\hat{\pi}^{l}_{\text{\tt RS}}\left(i\right)}-M_{i,\pi^{*}(i)}\right)$
$\leq\varepsilon L+2m2^{-L}\|M\|_{\max}$
$\leq\varepsilon L+2\varepsilon\|M\|_{\max}$
@@ Appendix G Proof of results from Section 8: Regret Minimization
In this section, we present the regret analysis of Algorithm 2. We postpone the proof of Theorem 8.1 to Section G.2, and first repeat the analysis of SupLinUCB algorithm given in (Takemura et al., 2021), culminating in Theorem G.1, which we then use in proving our upper bound on regret of RS-RMIN in Section G.2.
@@ G.1 Misspecified linear bandits
The reward of action $a$ in round $t$ is expressed as $r_{t,a}=x_{t,a}^{\top}\theta+\epsilon_{t,a}+\xi_{t}$. We assume that the misspecification is bounded, i.e., $\sup_{t,a}|\epsilon_{t,a}|\leq\epsilon_{\max}$ and that $\max_{t,a}|r_{t,a}|=\Theta(1)$. The noise process $\{\xi_{t}\}_{t\geq 1}$ is i.i.d. with $1$-subgaussian distribution. The regret up to round $T$ of an algorithm $\pi$ selecting $a_{t}^{\pi}$ in round $t$ is defined by:
@@
where $a_{t}^{\star}=\operatorname*{arg\,max}_{a\in[A]}\mathbb{E}[r_{t,a}]$. To solve the aforementioned misspecified linear bandits, we apply Algorithm 1 in (Takemura et al., 2021), a variant of SupLinUCB (Chu et al., 2011). The pseudo-code is presented as Algorithm 5.
@@
Let $L>0$ be such that $\|\theta\|_{2}\leq L$, and let $\Lambda=(1/L^{2})I_{d}$. Moreover, let $\delta\in(0,1)$, $T>0$ and define the threshold
$\beta(\delta)=\sqrt{2\log\left(\frac{TK}{\delta}\left\lceil\frac{1}{2}\log\frac{T}{d}\right\rceil\right)}+1.$ (44)
@@
$\mathcal{R}^{\pi}(T)\leq C\left(\beta(\delta)\sqrt{dT}\!+\epsilon_{\max}\sqrt{d}T+d\right)\log(T+1)\log(1+L^{4}T/d),$
@@
The result of Theorem G.1 entails that Algorithm 5 has an expected regret [^14] The regret guarantee of Theorem G.1 is stated with high probability, but the result can be immediately extended to a regret in expectation by choosing for example $\delta=1/T$. of order $\widetilde{O}(\sqrt{dT\log(K)}+\epsilon_{\max}\sqrt{d}T)$, where $\widetilde{O}$ hides polylogarithmic factors in $d$ and $T$. Therefore, SupLinUCB enjoys tight dependence in the dimension in both the term $\widetilde{O}(\epsilon_{\max}\sqrt{d}T)$ that is due to the misspecification (Lattimore et al., 2020) and in the term $\sqrt{dT\log(K)}$ that corresponds to the minimal regret without misspecification (Auer, 2002; Chu et al., 2011).
@@
Let us denote by $j_{t}^{\star}$ the optimal arm when given context $i_{t}$, i.e., let $j_{t}^{\star}=\pi^{\star}(i_{t})$ for all $t\in[T]$. We split the regret into the sum of two terms corresponding to the uniform exploration phase and a second phase where we apply Algorithm 2:
$R^{\pi}(T)=\underbrace{\sum_{t=1}^{T_{1}}\mathbb{E}[M_{i_{t},j^{\star}_{t}}-M_{i_{t},j_{t}^{\pi}}]}_{\begin{subarray}{c}R_{1}^{\pi}\\ \textit{(Regret of phase 1)}\end{subarray}}+\underbrace{\sum_{t=T_{1}+1}^{T}\mathbb{E}[M_{i_{t},j^{\star}_{t}}-M_{i_{t},j_{t}^{\pi}}]}_{\begin{subarray}{c}R_{2}^{\pi}\\ \textit{(Regret of phase 2)}\end{subarray}}$
Step 1: (Regret of phase 1). The first phase lasting for $T_{1}$ rounds accumulates at most $2T_{1}\|M\|_{\max}$ regret since $|M_{i,j}|\leq\|M\|_{\max}$ for all $(i,j)\in[m]\times[n]$. More precisely, we have
$R_{1}^{\pi}\leq 2T_{1}\|M\|_{\max}\leq C_{1}T_{1}.$ (45)
where we used that $\|M\|_{\max}=\Theta(1)$.
Step 2: (Regret of phase 2). In the second phase, which lasts for $T_{2}=T-T_{1}$ rounds, we can identify our observations $M_{i_{t},j_{t}}+\xi_{t}=\phi_{i_{t},j_{t}}^{\top}\theta+\epsilon_{i_{t},j_{t}}+\xi_{t}$ with rewards $r_{t,j_{t}}$ from the previous section, where we define the set of arms for the context $i_{t}$ by setting $(x_{t,a})_{a\in[n]}=(\phi(i_{t},j))_{j\in[n]}$ with $K=n$, $x_{a^{\star}_{t}}=\phi(i_{t},j^{\star}_{t})$ and $d=r(m+n)-r^{2}$. Thus, the regret $R_{2}^{\pi}$ generated during this phase can be equivalently written as:
$R_{2}^{\pi}=\sum_{t=T_{1}+1}^{T}\mathbb{E}[r_{t,a^{\star}_{t}}-r_{t,a_{t}^{\pi}}]=\mathbb{E}[{\cal R}^{\pi}(T_{2})]$
@@
${\cal R}^{\pi}(T_{2})\leq C_{2}\sqrt{r(m+n)}\left(\sqrt{T\log\left(\frac{Tmn}{\delta}\right)}+\epsilon_{\max}T+\sqrt{r(m+n)}\right)\log^{2}(1+mnT)$ (46)
where we used that $T_{2}\leq T$. To bound the second term in the parenthesis, we use the upper bound on the misspecification $\epsilon$ from Proposition 5.1. Indeed, we have that with probability at least $1-\delta$:
$\epsilon_{\max}=\max_{i,j}|\epsilon_{i,j}|\leq\frac{C\mu^{6}\kappa^{4}r^{3}(m+n)}{T_{1}\omega_{\min}\sqrt{mn}(m\wedge n)}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right),$ (47)
provided that $T_{1}\geq\frac{c\mu^{4}\kappa^{2}r^{2}(m+n)}{\omega_{\min}mn}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$, for some universal constants $C,c>0$. Define the event $\mathcal{E}$ as the event for which (47) holds. Next, we decompose $R_{2}^{\pi}$ as follows:

### v2

@@ D.3 Additional Lemmas
The control of the lower bound identified above is tight when $\omega$ is homogeneous and $\kappa=\Theta\left(1\right)$, since in this case, the two sides of the inequality only differ by a constant factor. In particular, the relaxed lower bound then scales as the lower bound of Theorem 6.3.
@@
We are now ready to derive the worst-case lower bound of Proposition 6.4. To that end, we will show that
$K_{M,\pi}:=\frac{1}{\sigma_{1}\omega_{\max}}\max\left(\sum_{i}\|Q^{i}_{\pi}\|_{2}^{2},\sum_{j}\|P^{j}_{\pi}\|_{2}^{2}\right)$
is of order $\frac{m+n}{\omega_{\max}mn}$ in the worst case (i.e., for a choice of homogeneous reward matrix and target policy that maximizes the lower bound). We use the notation $\pi_{j}:=\sum_{i=1}^{m}\rho_{i}\pi\left(j|i\right)$. We note that the third bullet point of the following proposition corresponds to the statement of Proposition 6.4.
@@
    $T\gtrsim\frac{\sigma^{2}\left(m+n\right)}{\omega_{\max}mn\varepsilon^{2}}\operatorname{kl}\left(\delta,1-\delta\right).$
@@
Since $P=U\sqrt{\Sigma}$ and $Q=V\sqrt{\Sigma}$, we have $\max_{i}\|P_{i}\|_{2}^{2}=\max_{i}\sum_{j=1}^{r}\sigma_{j}U_{x,j}^{2}\leq\sigma_{1}\|U\|_{2\to\infty}^{2}$ and similarly, $\max_{j}\|Q_{j}\|_{2}^{2}\leq\sigma_{1}\|V\|_{2\to\infty}^{2}.$ This finally ensures that
@@
$T\geq\frac{\sigma^{2}\operatorname{kl}\left(\delta,1-\delta\right)}{2\left(m\vee n\right)\omega_{\max}\varepsilon^{2}}.$
@@ Appendix F Proofs of results from Section 7: Best Policy Identification
In this appendix, we denote by $\Pi$ the set of deterministic policies (in other words, the set of functions from $[m]$ to $[n]$). To ease the notation, we also denote by $\hat{v}^{\pi}_{\text{\tt RS}}$ the RS-PE estimator of $v^{\pi}$ for a given policy $\pi$, and by $\hat{\pi}_{\text{\tt RS}}$ the policy learned by RS-BPI.
@@
$T\gtrsim\frac{\sigma^{2}\mu^{2}r\left(m+n\right)}{\omega_{\min}mn\,\varepsilon^{2}}\log\left(\frac{m+n}{\delta}\right)+\frac{K_{0}}{\varepsilon}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$
where $K_{0}$ has been defined in (28).
@@
By replacing $\phi_{\pi}$ by $\phi_{i,j}$ in the error analysis performed in Appendix D, we directly obtain that for every (context, arm) pair $(i,j)$, with probability larger than $1-\frac{\delta}{mn}$,
@@
$T\gtrsim\frac{\sigma^{2}\mu^{2}r\left(m+n\right)}{\omega_{\min}mn\,\varepsilon^{2}}\log\left(\frac{emn}{\delta}\right)+\frac{K_{0}}{\varepsilon}\log^{3}\left(\frac{mn(m+n)T}{\delta}\right).$
@@
with probability larger than $1-\delta,$ where $K_{0}$ has been defined in (28).
@@
$v^{*}-v^{\hat{\pi}_{\text{\tt RS}}}\leq\varepsilon\text{ for }T\gtrsim\frac{\sigma^{2}\mu^{2}r\left(m+n\right)}{\varepsilon^{2}}\log\left(\frac{m+n}{\delta}\right)+\frac{K_{0}}{\varepsilon}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$
for
$K_{0}\lesssim\frac{L^{2}\mu^{6}\kappa^{4}r^{3}\left(m+n\right)^{2}\sqrt{mn}}{\|M\|_{\max}\left(m\wedge n\right)}.$ (44)
We finally recall that from Remark D.8, $K_{0}\lesssim\sigma mn$ when $M$ is also homogeneous and $\|M\|_{\max}=\Theta(\sigma).$
We can of course derive the corresponding result for SBPI in the same way. More specifically, we obtain that $v^{\star}-v^{\hat{\pi}_{\text{\tt SBPI}}}\leq\varepsilon$ with probability larger than $1-\delta$ as soon as
$T\gtrsim\frac{L^{2}\mu^{6}\kappa^{4}r^{3}(m+n)mn}{(m\wedge n)^{2}\varepsilon^{2}}\log^{3}\!\left(\frac{(m+n)T}{\delta}\right)\!$
by combining the max-norm guarantee on $\widehat{M}$ of Proposition D.1 with the previous arguments171717to get the exact statement of the theorem, note that $\left(m+n\right)mn<\left(m+n\right)^{3}$.. ∎
@@ F.2 General context distribution case
We now explain how to drop the homogeneous context assumption in Theorem 7.1 while retaining the optimal scaling up to logarithmic factors with the technique highlighted in (Lee et al., 2023). The idea is to split the set of contexts into subsets on which the context distribution is homogeneous and to apply RS-BPI or SBPI on each subset. Let
$\mathcal{L}=\left\lceil\log_{2}\left(m/\varepsilon\right)\right\rceil$ (45)
and define $\mathcal{C}_{\mathcal{L}}=\left\{i\in[m],2^{-l-1}<\rho_{i}\leq 2^{-l}\right\}$ for $10\leq l\leq\mathcal{L}-$ 1, and $\mathcal{C}_{\mathcal{L}}=\left\{i\in[m],\rho_{i}\leq 2^{-\mathcal{L}}\right\}$. On each $\mathcal{C}_{l}$ for $10\leq l\leq\mathcal{L}-$ 1, the distribution defined by the normalized context weights is homogeneous. On $\mathcal{C}_{\mathcal{L}}$, the context weights are small enough to not meaningfully contribute to the BPI error. We summarize below the algorithm structure for RS-BPI, and we recall that $\alpha$ is the proportion of samples used in the first phase of the algorithm. Of course, a SBPI version of this algorithm can be constructed in the same way.
@@
  Input: Budget of rounds $T$, data splitting parameter $\alpha$, context distribution $\rho$, regularization parameter $\tau$
@@
  Output: $\hat{\pi}_{\text{\tt RS}}$ such that $\hat{\pi}_{\text{\tt RS}}=\hat{\pi}^{l}_{\text{\tt RS}}$ on $\mathcal{C}_{l}$ for $10\leq l\leq\mathcal{L}-$ 1 and arbitrary on $\mathcal{C}_{\mathcal{L}}$.
@@
Let $\varepsilon>0$, $\delta\in\left(0,1\right)$, and define $\rho\left(\mathcal{C}_{l}\right):=\sum_{i\in\mathcal{C}_{l}}\rho_{i}$. Assume that $\varepsilon<\left(\displaystyle\min_{0\leq l\leq\mathcal{L}-1}\rho\left(\mathcal{C}_{l}\right)\right)\|M\|_{\max}$. With the choices $T_{1}=\lfloor\alpha T\rfloor$ and $\tau\leq r(m\wedge n)^{-1}(\rho_{\min}/\rho_{\max})\log(16d/\delta)$, $v^{\star}-v^{\hat{\pi}_{\text{\tt RS}}}\leq\varepsilon^{\prime}$ with probability larger than $1-\delta$ as soon as
$T\gtrsim\frac{\sigma^{2}\mu^{2}r\left(m+n\right)}{\varepsilon^{2}}\log\left(\frac{\mathcal{L}\left(m+n\right)}{\delta}\right)+\frac{K_{0}}{\varepsilon}\log^{3}\left(\frac{\mathcal{L}(m+n)T}{\delta}\right)\!,$
where $K_{0}$ has been defined in (44), $\mathcal{L}$ has been defined in (45) and $\varepsilon^{\prime}=\varepsilon\left(\mathcal{L}+2\|M\|_{\max}\right)$.
@@
Let $10\leq l\leq\mathcal{L}-$ 1 and $\delta_{l}=\delta/\mathcal{L}$. From Lemma 9 in (Lee et al., 2023), we know that with high probability, at least
$\frac{T\rho\left(\mathcal{C}_{l}\right)}{2}\gtrsim\frac{\sigma^{2}\mu^{2}r\left(m+n\right)}{\varepsilon_{l}^{2}}\log\left(\frac{m+n}{\delta_{l}}\right)+\frac{K_{0}}{\varepsilon_{l}}\log^{3}\left(\frac{(m+n)T}{\delta_{l}}\right)$
out of the $T$ samples correspond to a context in $\mathcal{C}_{l}$ for $\varepsilon_{l}:=\varepsilon/\rho\left(\mathcal{C}_{l}\right).$ Since $\varepsilon_{l}\leq\|M\|_{\max}$, by Theorem 7.1, Algorithm 1 on $\mathcal{C}_{l}$ yields a policy $\hat{\pi}^{l}_{\text{\tt RS}}$ such that
$\sum_{i\in\mathcal{C}_{l}}\frac{\rho_{i}}{\rho\left(\mathcal{C}_{l}\right)}\left(M_{i,\hat{\pi}^{l}_{\text{\tt RS}}(i)}-M_{i,\pi^{\star}(i)}\right)\leq\frac{\varepsilon}{\rho\left(\mathcal{C}_{l}\right)}$
@@
$\sum_{i=1}^{m}\rho_{i}\left(M_{i,\hat{\pi}_{\text{\tt RS}}(i)}-M_{i,\pi^{\star}(i)}\right)$ $=\sum_{l=0}^{\mathcal{L}-1}\sum_{i\in\mathcal{C}_{l}}\rho_{i}\left(M_{i,\hat{\pi}^{l}_{\text{\tt RS}}(i)}-M_{i,\pi^{\star}(i)}\right)+\sum_{i\in\mathcal{C}_{\mathcal{L}}}\rho_{i}\left(M_{i,\hat{\pi}^{l}_{\text{\tt RS}}\left(i\right)}-M_{i,\pi^{\star}(i)}\right)$
$\leq\varepsilon\mathcal{L}+2m2^{-\mathcal{L}}\|M\|_{\max}$
$\leq\varepsilon\mathcal{L}+2\varepsilon\|M\|_{\max}$
@@ Appendix G Proof of results from Section 8: Regret Minimization
In this appendix, we present the regret analysis of Algorithm 2. We postpone the proof of Theorem 8.1 to Section G.2, and first repeat the analysis of SupLinUCB algorithm given in (Takemura et al., 2021), culminating in Theorem G.1, which we then use to prove our upper bound on the regret of RS-RMIN in Section G.2.
@@ G.1 Misspecified linear bandits
The reward of action $a$ in round $t$ is expressed as $r_{t,a}=x_{t,a}^{\top}\theta+\epsilon_{t,a}+\xi_{t}$. We assume that the misspecification is bounded, i.e., $\sup_{t,a}|\epsilon_{t,a}|\leq\epsilon_{\max}$ and that $\max_{t,a}|r_{t,a}|=r_{\max}$. The noise process $\{\xi_{t}\}_{t\geq 1}$ is i.i.d. with $\sigma$-subgaussian distribution. The regret up to round $T$ of an algorithm $\pi$ selecting $a_{t}^{\pi}$ in round $t$ is defined by:
@@
where $a_{t}^{\star}=\operatorname*{arg\,max}_{a\in[A]}\mathbb{E}[r_{t,a}]$. To solve the misspecified linear bandit problem, we apply Algorithm 1 in (Takemura et al., 2021), a variant of SupLinUCB (Chu et al., 2011). The pseudo-code is presented in Algorithm 5.
@@
Let $B>0$ be such that $\|\theta\|_{2}\leq B$, and let $\Lambda=(\sigma^{2}/B^{2})I_{d}$. Moreover, let $\delta\in(0,1)$, $T>0$ and define the threshold
$\beta(\delta)=\sigma(1+\sqrt{2\log\left(\frac{TK}{\delta}\left\lceil\frac{1}{2}\log\frac{T}{d}\right\rceil\right)}).$ (46)
@@
$\mathcal{R}^{\pi}(T)\leq C\left(\sigma\sqrt{dT\log K}\!+\epsilon_{\max}\sqrt{d}T+dr_{\max}\right)\log(T+1)\log\left(1+\frac{TB^{2}}{d\sigma^{2}}\right),$
@@
The result of Theorem G.1 entails that Algorithm 5 has an expected regret [^18] The regret guarantee of Theorem G.1 is stated with high probability, but the result can be immediately used to bound the regret in expectation by choosing for example $\delta=1/T$. of order $\widetilde{O}(\sqrt{dT\log(K)}+\epsilon_{\max}\sqrt{d}T)$, where $\widetilde{O}$ hides polylogarithmic factors in $d$ and $T$. Therefore, SupLinUCB enjoys tight dependence in the dimension in both the term $\widetilde{O}(\epsilon_{\max}\sqrt{d}T)$ that is due to the misspecification (Lattimore et al., 2020) and in the term $\sqrt{dT\log(K)}$ that corresponds to the minimal regret without misspecification (Auer, 2002; Chu et al., 2011).
@@
Let us denote by $j_{t}^{\star}$ the optimal arm given context $i_{t}$, i.e., let $j_{t}^{\star}=\pi^{\star}(i_{t})$ for all $t\in[T]$. We split the regret into the sum of two terms corresponding to the uniform exploration phase and a second phase where we apply Algorithm 2:
$R^{\pi}(T)=\underbrace{\sum_{t=1}^{T_{1}}\mathbb{E}[M_{i_{t},j^{\star}_{t}}-M_{i_{t},j_{t}^{\pi}}]}_{\begin{subarray}{c}R_{1}^{\pi}\\ \textit{(Regret of phase 1)}\end{subarray}}+\underbrace{\sum_{t=T_{1}+1}^{T}\mathbb{E}[M_{i_{t},j^{\star}_{t}}-M_{i_{t},j_{t}^{\pi}}]}_{\begin{subarray}{c}R_{2}^{\pi}\\ \textit{(Regret of phase 2)}\end{subarray}}.$
Step 1: (Regret of phase 1). Since $|M_{i,j}|\leq\|M\|_{\max}$ for all $(i,j)\in[m]\times[n]$, the first phase lasting for $T_{1}$ rounds accumulate a regret of at most:
$R_{1}^{\pi}\leq 2T_{1}\|M\|_{\max}.$ (47)
Step 2: (Regret of phase 2). In the second phase, which lasts for $T_{2}=T-T_{1}$ rounds, we can identify our observations $M_{i_{t},j_{t}}+\xi_{t}=\phi_{i_{t},j_{t}}^{\top}\theta+\epsilon_{i_{t},j_{t}}+\xi_{t}$ with rewards $r_{t,j_{t}}$ from the previous section, where we define the set of arms for the context $i_{t}$ by setting $(x_{t,a})_{a\in[K]}=(\phi(i_{t},j))_{j\in[n]}$ with $K=n$, $x_{a^{\star}_{t}}=\phi(i_{t},j^{\star}_{t})$ and $d=r(m+n)-r^{2}$. Thus, the regret $R_{2}^{\pi}$ generated during this phase can be equivalently written as:
$R_{2}^{\pi}=\sum_{t=T_{1}+1}^{T}\mathbb{E}[r_{t,a^{\star}_{t}}-r_{t,a_{t}^{\pi}}]=\mathbb{E}[{\cal R}^{\pi}(T_{2})].$
@@
${\cal R}^{\pi}(T_{2})\leq C_{2}\sqrt{r(m+n)}\left(\sigma\sqrt{T\log\left(\frac{Tmn}{\delta}\right)}+\epsilon_{\max}T+\|M\|_{\max}\sqrt{r(m+n)}\right)\log^{2}\left(T(m+n)\frac{L^{2}}{\sigma^{2}}\right)$ (48)
where we used that $T_{2}\leq T$. To bound the second term in the parenthesis, we use the upper bound on the misspecification $\epsilon_{\max}$ of Corollary 5.1. Indeed, we have that with probability at least $1-\delta$:
$\epsilon_{\max}\leq\frac{CL^{2}\kappa^{3}\mu^{2}r(m+n)}{\sigma_{r}T_{1}\omega_{\min}(m\wedge n)}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right),$ (49)
provided that $T_{1}\geq c\frac{L^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$, for some universal constants $C,c>0$. Define $\mathcal{E}$ as the event under which (49) holds. Next, we decompose $R_{2}^{\pi}$ as follows: